# Towards a Theory of Pragmatic Information

### v5

@@ Abstract
Standard information theory says nothing about how much meaning is conveyed by a message. We fill this gap with a rigorously justifiable, quantitative definition of “pragmatic information”, the amount of meaning in a message relevant to a particular decision. We posit that such a message updates a random variable, $\omega$, that informs the decision. The pragmatic information of a single message is then defined as the Kulbach-Leibler divergence between the apriori and aposteriori probabilities of $\omega$; the pragmatic information of a message ensemble is the expected value of the pragmatic information of the ensemble’s component messages. We justify these definitions by proving that the pragmatic information of a single message is the expected difference between the shortest binary encoding of $\omega$ under the a priori and a posteriori distributions, and that the average of the pragmatic values of individual messages, when sampled a large number of times from the ensemble, approaches its expected value.
@@ 1 Introduction
Although the primary goal of the present work is to advance the theory as such, any theory that is termed “pragmatic” should have practical applications, so we mention a few. [3] proposed that the rate at which biological evolution aggregates “meaningful information” could be used to measure the rate at which evolution proceeds. Throughout the present paper, we present a second application, namely the role that the history of previous payouts of a slot machine acquire meaning in the context of deciding whether to keep playing. We also show how our conceptual framework makes possible a reformulation of the celebrated efficient market hypothesis of financial economics as the claim that none of the available information about a given security is “pragmatically useful”.
@@ 2.1 Our Conceptual Framework
While our conceptual framework can be applied to the wealth of so-called “multi-armed bandit problems” (See, for example, [11] and [12].), we consider the following, even simpler “one-armed bandit” problem: we can pay \$1 to play a slot machine (also known as a “one armed bandit”) as many times as we wish. After each play, we will win a \$2 payout with some constant probability $\pi$, or nothing, with probability $1-\pi$, with the payout events independent from one play to the next. The problem is that $\pi$ is not known in advance!
Suppose we denote the result of the $t^{th}$ such play by the random variable $\mathbf{1}_{t}$, with
$\mathbf{1}_{t}=\begin{cases}1&\quad\text{if there is a payout on the $t^{th}$ play}\\ 0&\quad\text{otherwise}\end{cases}$
@@
per play remains constant from one play to the next, so that the Shannon information for the $T$ independent plays is $T[\pi\log_{2}\pi+(1-\pi)\log_{2}(1-\pi)]$. However, the Shannon information captures data such as the particular order of successes and failures, which is more than we need to know to estimate $\pi$.
@@ 2.2 Defining the Pragmatic Information of a Single Message
The formula that [9] proposed for “surprise” is essentially the same as (2) for a single message (Below, we will generalize this definition to any $m$ in some ensemble of possible messages.). However, [9] left ambiguous whether to define this quantity as (in our notation) $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ or $D_{\Delta}(\mathbf{q}||\mathbf{p}_{m})$. They prefer the latter, as it easier for them to compute intuition-building analytic results in specific cases; however, we prefer the former, as we assume – perhaps naively! – that $\mathbf{p}_{m}$ really does provide an improvement over $\mathbf{q}$. Our preferred definition also makes more sense in subsequent sections of this paper, where we list some known/easily derivable properties [^4] See [15] for proofs of all of the theorems in the next subsection, except for our “Wrong Code Theorem”, which is an embellishment of their results, and the obvious corollary to Theorem 2.3.4. of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ that make it an appealing choice for a measure of pragmatic information of a single message.
@@ 2.3 Properties of the Pragmatic Information of a Single Message
${\mathcal{H}}(\mathbf{p}_{m})+D(\mathbf{p}_{m}||\mathbf{q})\leq{\mathcal{E}}\left[L_{\mathbf{q}}(\omega)\right]]<{\mathcal{H}}(\mathbf{p}_{m})+D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})+1,$
@@
We conclude that $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ can reasonably be interpreted as the amount of information that $\Delta$ has “learned” from $m$.
@@
In the special case where $\Delta$ and $\Delta^{\prime}$ are completely independent of each other, i.e. that $\{q^{\prime}_{i^{\prime}|i}\}=\mathbf{q}^{\prime}$ and $\{p^{\prime}_{i^{\prime}|i,m,m^{\prime}}\}=\mathbf{p}^{\prime}_{m}$,
@@
In the special case where $\mathbf{q}=\mathbf{q}^{\prime}=\mathbf{q}_{\lambda}$ and ${\mathbf{u}}_{j}$ and ${\mathbf{u}}_{k}$ are unit vectors in the $j^{\rm th}$ and $k^{\rm th}$ direction, respectively,
$D\Big{(}\lambda{\mathbf{u}}_{j}+(1-\lambda){\mathbf{u}}_{k}||\mathbf{q}\Big{)}$ $\leq\lambda D\left({\mathbf{u}}_{j}||\mathbf{q}\right)+(1-\lambda)D\left({\mathbf{u}}_{k}||\mathbf{q}\right)$
$\leq\max\left(-\log q_{j},-\log q_{k}\right).$
It follows that, for given $\mathbf{q}$, we have the following
@@
The maximum value of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ is $\max_{l}(-\log q_{l})$, with the max taken over all of the $q$’s for which $p_{l|m}>0$.
In other words, the pragmatic information conveyed by a single message $m$ is a maximum when $\Delta$ treats the outcome indicated by $m$ as a certainty.
@@ 3 The Pragmatic Information of an Ensemble of Messages
with probability 1, with the subscript being omitted from the expectation because the distribution of the $M_{k}$’s is the same for all $k$. We stress that, as in the previous section, $\Delta$ makes each of its decisions as though none of the previous decisions had ever been made. However, the hypotheses of the ergodic theorem are sufficiently weak that no such assumption is required for the $M_{k}$’s. While they could be sampled independently, they could also be a realization of a Markov chain, etc. The generality of this result therefore justifies the following
@@ 3.1 An Example: The One Armed Bandit Problem
$=p\left\{\frac{w+2}{T+3}\log\left[\frac{(w+2)(T+2)}{(T+3)(w+1)}\right]+\frac{T-w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]\right\}$
$\qquad+(1-\pi)\left\{\frac{w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]+\frac{T-w+2}{T+3}\log\left[\frac{(T-w+2)(T+2)}{(T+3)(T-w+1)}\right]\right\}$
As $T$ gets large, $w=\pi T+o(T)$ with probability 1, by the Kolmogorov Strong Law of Large Numbers [17]. Straightforward algebraic manipulation then shows that the arguments of all of the logarithms in $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ approach 1 monotonically as $T$ increases. We conclude that $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ steadily decreases to zero with increasing $T$.
@@ 3.2 Properties of the Pragmatic Information of an Ensemble of Messages
The theorems in the previous section remain true for ${\Phi_{\Delta}(\mathcal{M};\Omega)}$, since it is just a weighted sum of Kullbach-Leibler divergences. We therefore have immediately
@@
$\leq-\log q_{l}$
where, for each $m$, $k(m)=\operatorname*{arg\,max}(-\log q_{l})$, with the max taken over all of the $q$’s for which $p_{l|m}>0$ and $-\log q_{l}$ is the largest of the $-\log q$’s (The $\varphi$’s sum to 1.).
@@
  • The joint probability that these messages are received and the decision makers output $\omega_{i}$ and $\omega^{\prime}_{i^{\prime}}$, respectively, is $p_{i,i^{\prime},m,m^{\prime}}$.
@@
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})$ $=E\left[D(\{p_{i,i^{\prime}|M,M^{\prime}}\}||\{q_{i,i^{\prime}}\})\right]$
@@
As an example, suppose $\Delta$ and $\Delta^{\prime}$ have the respective Boolean state of the world variables $\omega$ and $\omega^{\prime}$, informed by Boolean messages $M$ and $M^{\prime}$ according to the formulas $\omega=M$ and $\omega^{\prime}=M\ \mathtt{AND}\ M^{\prime}$ with probability one. Assume further that $M$ and $M^{\prime}$ are independent random variables, each assuming the values 1 and 0 with equal probability. Thus, a priori, $\omega$ also assumes the values 1 and 0 with equal probability, $\omega^{\prime}$ is 1 with probability ${1\over 4}$ and 0 with probability ${3\over 4}$, and $p_{i,i^{\prime}|m,m^{\prime}}=p_{i^{\prime}|i,m,m^{\prime}}=1$ for $(i,i^{\prime},m,m^{\prime})=(0,0,0,0)$, $(i,i^{\prime},m,m^{\prime})=(0,0,0,1)$, $(i,i^{\prime},m,m^{\prime})=(1,0,1,0)$, and $(i,i^{\prime},m,m^{\prime})=(1,1,1,1)$ and zero otherwise. We then have
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M};\Omega,\Omega^{\prime})={3\over 2},$
with
$\Phi_{\Delta}({\mathcal{M}};\Omega)=1,$
and
$\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)={1\over 2}.$
We also note that conditioning need not decrease pragmatic information, as it does with Shannon entropy, since
$\Phi_{\Delta^{\prime}}({\mathcal{M}^{\prime}};\Omega^{\prime})$ $={3\over 2}-{3\over 4}\log 3\approx 0.3113$
$<{1\over 2}=\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)$
In this case, the receipt of $M$ has provided a context in which $M^{\prime}$ has become more informative.

### v6

@@ Abstract
Standard information theory says nothing about how much meaning is conveyed by a message. We fill this gap with a rigorously justifiable, quantitative definition of “pragmatic information”, the amount of meaning in a message relevant to a particular decision. We posit that such a message updates a random variable, $\omega$, that informs the decision. The pragmatic information of a single message is then defined as the Kulbach-Leibler divergence between the prior and posterior probabilities of $\omega$; the pragmatic information of a message ensemble is the expected value of the pragmatic information of the ensemble’s component messages. We justify these definitions by proving that the pragmatic information of a single message is the expected difference between the shortest binary encoding of $\omega$ under the a priori and a posteriori distributions, and that the average of the pragmatic values of individual messages, when sampled a large number of times from the ensemble, approaches its expected value.
@@ 1 Introduction
Although the primary goal of the present work is to advance the theory as such, any theory that is termed “pragmatic” should have practical applications, so we mention a few. [3] proposed that the rate at which biological evolution aggregates “meaningful information” could be used to measure the rate at which evolution proceeds. Throughout the present paper, we present a second application, namely how knowledge of previous payouts of a slot machine acquires meaning in the context of deciding whether to keep playing. A final application is a reformulation of the celebrated efficient market hypothesis of financial economics as the claim that none of the available information about a given security is “pragmatically useful”.
@@ 2.1 Our Conceptual Framework
While our conceptual framework can be applied to the wealth of so-called “multi-armed bandit problems” (see, for example, [9], [11] and [10].), we consider the following, even simpler problem: we have to decide whether to pay \$1 to play a slot machine (also known as a “one armed bandit”). We know that will win a \$2 payout with some probability $\pi$, or nothing, with probability $1-\pi$. We have access to the outcomes of as many previous plays as we wish. We also know that the payout probability is $\pi$ for each play, and each play is independent of the others. The problem is that we don’t know the value of $\pi$!
Suppose we denote the outcome of the $t^{th}$ such play by the random variable $\mathbf{1}_{t}$, with
$\mathbf{1}_{t}=\begin{cases}1&\quad\text{if we win on the $t^{th}$ play}\\ 0&\quad\text{otherwise}\end{cases}$
@@
per play remains constant from one play to the next. However, the Shannon information captures data such as whether particular plays are wins or losses, which is more than we need to know in order to estimate $\pi$ using (1).
@@ 2.2 Defining the Pragmatic Information of a Single Message
The formula that [7] proposed for “surprise” is essentially the same as (2) for a single message. Below, we will generalize this definition to any $m$ in some ensemble of possible messages. However, [7] left ambiguous whether to define this quantity as (in our notation) $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ or $D_{\Delta}(\mathbf{q}||\mathbf{p}_{m})$. They prefer the latter, as it easier for them to compute intuition-building analytic results in specific cases; however, we prefer the former, as we assume – perhaps naively! – that $\mathbf{p}_{m}$ really does provide an improvement over $\mathbf{q}$. Our preferred definition also makes more sense in subsequent sections of this paper, where we list some known/easily derivable properties [^4] See [13] for proofs of all of the theorems in the next subsection, except for our “Wrong Code Theorem”, which is an embellishment of their results, and the corollary to Theorem 2.3.4. of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ that make it an appealing choice for a measure of pragmatic information of a single message.
@@ 2.3 Properties of the Pragmatic Information of a Single Message
${\mathcal{H}}(\mathbf{p}_{m})+D(\mathbf{p}_{m}||\mathbf{q})\leq{\mathcal{E}}\left[L_{\mathbf{q}}(\omega)\right]<{\mathcal{H}}(\mathbf{p}_{m})+D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})+1,$
@@
We conclude that $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ can reasonably be interpreted as the amount of information that $\Delta$ has “learned” from $m$, assuming that $\mathbf{p}_{m}$ really is a better estimate of the true probablilty distribution of $\omega$ than $\mathbf{q}$.
@@
In the special case where $\Delta$ and $\Delta^{\prime}$ are completely independent of each other, i.e. that $\{q^{\prime}_{i^{\prime}|i}\}=\mathbf{q}^{\prime}$ and $\{p^{\prime}_{i^{\prime}|i,m,m^{\prime}}\}=\mathbf{p}^{\prime}_{m^{\prime}}$,
@@
It is useful at this point to introduce the following
Definition.
A message $m$ is pragmatically definitive for a given $\Delta$ if $\mathbf{p}_{m}=\mathbf{u}_{k}$ for some $k$, i.e. the unit vector in the $k^{\rm th}$ direction.
In other words, $\Delta$ is certain that outcome $k$ will be achieved if it receives a pragmatically definitive message to that effect. Note that all that pragmatic definitiveness guarantees is the definitive prediction of an outcome, but not necessarily the desired one!
Using this terminology, we have for given $\mathbf{q}$, the following
@@
The maximum value of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ is the maximum of all possible pragmatically definitive messages, namely $\max_{i}(-\log q_{i})$.
Proof.
Since $\mathbf{p}_{m}=\sum_{i}p_{i|m}\mathbf{u}_{i}$, where $\mathbf{u}_{i}$ is a unit vector in the ${\rm i}^{th}$ direction, it follows from repeated applications of Theorem 2.3.4 that
$D\Big{(}\mathbf{p}_{m}||\mathbf{q}\Big{)}$ $=D\Big{(}\sum_{i}p_{i|m}\mathbf{u}_{i}||\mathbf{q}\Big{)}$
$\leq\sum_{i}p_{i|m}D\left(\mathbf{u}_{i}||\mathbf{q}\right)$
$\leq\max_{i}\left(-\log q_{i}\right)$
@@ 3 The Pragmatic Information of an Ensemble of Messages
with probability 1, with the expectation taken under the distribution of the $M_{k}$’s. We stress that, as in the previous section, $\Delta$ makes each of its decisions as though none of the previous decisions had ever been made. However, the hypotheses of the ergodic theorem are sufficiently weak that no such assumption is required for the $M_{k}$’s. While they could be sampled independently, they could also be a realization of a Markov chain, etc. The generality of this result therefore justifies the following
@@ 3.1 An Example: The One Armed Bandit Problem
$=\pi\left\{\frac{w+2}{T+3}\log\left[\frac{(w+2)(T+2)}{(T+3)(w+1)}\right]+\frac{T-w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]\right\}$
$\qquad+(1-\pi)\left\{\frac{w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]+\frac{T-w+2}{T+3}\log\left[\frac{(T-w+2)(T+2)}{(T+3)(T-w+1)}\right]\right\}.$
As $T$ gets large, $w=\pi T+o(T)$ with probability 1, by the Kolmogorov Strong Law of Large Numbers [15]. Straightforward algebraic manipulation then shows that the arguments of all of the logarithms in $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ approach 1 monotonically as $T$ increases. We conclude that $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ steadily decreases to zero with increasing $T$. Figure 1 shows this decrease, with $w=\pi T$, its most likely value. It also shows that the pragmatic information accruing from each trial is larger as $\pi$ gets closer to 1/2, as we would expect: the play/no play decision is less obvious in that case, so knowing the outcome of each additional previous trial is all the more informative.
Figure 1: Pragmatic information for the “one armed bandit”
@@ 3.2 Properties of the Pragmatic Information of an Ensemble of Messages
Because ${\Phi_{\Delta}(\mathcal{M};\Omega)}$ is just a weighted sum of Kullbach-Leibler divergences, we have immediately
@@
The following definition will be useful in the discussion below:
Definition.
A message ensemble $\mathcal{M}$ is pragmatically definitive for a given $\Delta$ if for all $m\in\mathcal{M}$, $m$ is pragmatically definitive.
For such ensembles, we have $p_{i|m}=\delta_{i,k(m)}$, i.e. $p_{i|m}=1$ if $i=k(m)$, the outcome to which $m$ is mapped deterministically and zero otherwise. It follows that
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $=\sum_{i,m}\varphi_{m}p_{i|m}\log\left(\frac{p_{i|m}}{q_{i}}\right)$
$=\sum_{i,m}\varphi_{m}\delta_{i,k(m)}\log\left(\frac{\delta_{i,k(m)}}{q_{i}}\right)$
$=-\sum_{m}\varphi_{m}\log q_{k(m)}$
for such ensembles. Since the $\varphi$’s sum to 1, we also have
@@
$\leq\max_{i}\left(-\log q_{i}\right)$
@@
  • The joint probability that these messages are received and the decision maker’s output $\omega_{i}$ and $\omega^{\prime}_{i^{\prime}}$, respectively, is $p_{i,i^{\prime},m,m^{\prime}}$.
@@
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})$ $=E\left[D(\{p_{i,i^{\prime}|m,ms^{\prime}}\}||\{q_{i,i^{\prime}}\})\right]$
@@
As an example, suppose $\Delta$ and $\Delta^{\prime}$ have the respective Boolean state of the world variables $\omega$ and $\omega^{\prime}$, and, a priori, we know nothing about them. Thus, we take $q_{i}=q_{i^{\prime}}=1/2$ and $q_{i,i^{\prime}}=q_{i}q_{i^{\prime}}=1/4$ for all values of $i$ and $i^{\prime}$. Suppose further that our knowledge is updated by Boolean messages $M$ and $M^{\prime}$ according to the formulas $\omega=M$ and $\omega^{\prime}=M\ \cap\ M^{\prime}$ with probability one. Finally, suppose that $M$ and $M^{\prime}$ are independent random variables, with $M$ assuming the values 1 and 0 with equal probability, and $M^{\prime}$ assuming the value 1 with probability ${1\over 4}$ and 0 with probability ${3\over 4}$. It follows that $p_{i,i^{\prime}|m,m^{\prime}}=1$ for $(i,i^{\prime},m,m^{\prime})=(0,0,0,0)$, $(i,i^{\prime},m,m^{\prime})=(0,0,0,1)$, $(i,i^{\prime},m,m^{\prime})=(1,0,1,0)$, and $(i,i^{\prime},m,m^{\prime})=(1,1,1,1)$ and zero otherwise. We then have
$\Phi_{\Delta,\Delta^{\prime}}(\mathcal{M},\mathcal{M}^{\prime};\Omega,\Omega^{\prime})$ $=\sum_{i,i^{\prime},m,m^{\prime}}\varphi_{m,m^{\prime}}p_{i,i^{\prime}|m,m^{\prime}}\log\left(\frac{p_{i,i^{\prime}|m,m^{\prime}}}{q_{i,i^{\prime}}}\right)$
$=\left(\frac{1}{2}\right)\left(\frac{3}{4}\right)(1)\log\left[\frac{1}{\left(\frac{1}{2}\right)\left(\frac{1}{2}\right)}\right]\ \ \textrm{ (the $(i,i^{\prime},m,m^{\prime})=(0,0,0,0)$ term)}$)
$\hphantom{=\ }+\left(\frac{1}{2}\right)\left(\frac{1}{4}\right)(1)\log\left[\frac{1}{\left(\frac{1}{2}\right)\left(\frac{1}{2}\right)}\right]\ \ \textrm{ (the $(i,i^{\prime},m,m^{\prime})=(0,0,0,1)$ term)}$)
$\hphantom{=\ }+\left(\frac{1}{2}\right)\left(\frac{3}{4}\right)(1)\log\left[\frac{1}{\left(\frac{1}{2}\right)\left(\frac{1}{2}\right)}\right]\ \ \textrm{ (the $(i,i^{\prime},m,m^{\prime})=(1,0,1,0)$ term)}$)
$\hphantom{=\ }+\left(\frac{1}{2}\right)\left(\frac{1}{4}\right)(1)\log\left[\frac{1}{\left(\frac{1}{2}\right)\left(\frac{1}{2}\right)}\right]\ \ \textrm{ (the $(i,i^{\prime},m,m^{\prime})=(1,1,1,1)$ term)}$)
$=2.$
In this example, $p_{i,i^{\prime}|m,m^{\prime}}=p_{i^{\prime}|i,m,m^{\prime}}$ for all values of $i,i^{\prime},m$ and $m^{\prime}$ because $m$ and $m^{\prime}$ completely determine $i^{\prime}$ without any need to know $i$. Also, $q_{i^{\prime}|i}=q_{i^{\prime}}$ because $q_{i}$ and $q_{i^{\prime}}$ are independent. $\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)$ can therefore be calculated exactly as above, except that the denominator of the quantity in the logarithm evaluates to 1/2 instead of the 1/4 that appears in the above. We conclude that $\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)=1$ in this example.
A similar, albeit simpler calculation shows that $\Phi_{\Delta}(\mathcal{M};\Omega)=1$ in the example, as must be the case, given the above chain rule. More surprising is the result that
$\Phi_{\Delta^{\prime}}(\mathcal{M}^{\prime};\Omega^{\prime})$ $=3/4$
$<\Phi_{\Delta^{\prime}|\Delta}(\mathcal{M}^{\prime};\Omega^{\prime}|\mathcal{M},\Omega)$
$=1.$
In other words, conditioning need not reduce pragmatic information, as it does with Shannon entropy, because the receipt of $M$ can provide a context in which $M^{\prime}$ has become more informative.